OpenStudy (anonymous):

cube root of 35/cube root of 28

OpenStudy (unklerhaukus):

\[=\frac{\sqrt[3]{35}}{\sqrt[3]28}=\sqrt[3]\frac{35}{28}=\sqrt[3] \frac{5} {4}\]